OpenStudy (sidmanfu):

Show algebraically that the vertices A (13,3) B (13,5) and C (=2.20) form a right angle triangle.

OpenStudy (anonymous):

1st off , you need to calculate , [AB] , [AC], [BC] we don't where the menu corner is , do you have anyidea what to do next ?

OpenStudy (sidmanfu):

is this to help find the slope of the triangle?

OpenStudy (sidmanfu):

nope not at all

OpenStudy (sidmanfu):

what do you mean by that because I'm already calculating the slope.

OpenStudy (anonymous):

we know that , AB² + AC² = BC² if BC is the longest , so A is going to be our square

OpenStudy (sidmanfu):

ohhhh so I figuerd out the slope for AB and AC now I need figure out BC

OpenStudy (sidmanfu):

AB's slope is \[\frac{ 2 }{ 16 }\]

OpenStudy (sidmanfu):

AC's slope is \[\frac{ -23 }{ -15 }\]

OpenStudy (anonymous):

how did you find those ?